If the area of the rectangle is 600 cm^2, what is the perimeter?
Please HINT me and give me TIPS on solving it, though, I would like to know how to do it myself, without the answer just given to me. Thank you.

Answer by John10(297)

HINT: By the extra info, the width is 2/3 the length
W = (2/3)L
A = W*L = 600
Solve for W and L then apply the formula Perimeter = 2(W + L)
John10:)